The results of a one-sample t test were t (18) = 2.11, p < 0.05. In this example, the sample size is (report as a whole number):
Answer #1

Given that, for one-tailed test, t(18) = 2.11

and p-value is less than 0.05 ( i.e. p < 0.05)

Here, degrees of freedom = n - 1 = 18

So, n = 18 + 1 = 19

Therefore, required sample size is 19